32.46 ml of nitrogen measured at STP is relesed when 0.1g of a primary amine is acted upon by excess of nitrous acid..       R-NH2+HNO2 gives R-OH+N2+H2O.calculate molar mass of amine..

Grade:11

1 Answers

Arun
3 years ago
Dear student
 
 
let its molar mass be m then by stiochiometry 0.1/m = 32.46/(1000*22.4) or m = 69g
 
Hope it helps